OpenStudy (anonymous):

A sporting goods store has added a section with trail foods and snacks. Raisins cost $5.00 per pound and peanuts cost $3.80 per pound. Raisins and peanuts are going to be sold in a mixture of 30 pounds which will cost $4.05 per pound? How many pounds of raisins will be in the mix? A. 6.1 lbs B. 6.25 lbs C. 5.5 lbs D. 5.75 lbs

hero (hero):

Hint: 5x + 3.80y = 4.05*30 x + y = 30